Three 3.0μC charges are at the three corners of a square of side 0.50m. The last corner is occupied by a -3.0μC charge. Find the electric field at the center of the square.

1
Expert's answer
2022-03-17T09:47:30-0400

Electric field due to charge

"q_1" And "q_3"

"E_1=E_3=\frac{kq_1}{r^2}=\frac{kq_3}{r^2}"


"E_1=E_3=\frac{9\times10^9\times3\times10^{-6}}{0.25^2}=432000N/C"

At symmetry "E_1,E_3" Are equal and opposite direction

Now net electric field due to charge

"q_2,q_4"

"E_{net}=E_2+E_4"


"E_{net}=\frac{9\times10^9\times3\times10^{-6}}{0.25^2}+\frac{9\times10^9\times3\times10^{-6}}{0.25^2}=432000+432000=864000N/C"